# JWST/MIRI-MRS view of the metal-poor galaxy CGCG 007-025: the spatial location of PAHs and very highly ionized gas

### v1

@@ 5.2 What PAH and line ratios reveal about CGCG 007-025
Complementing these diagnostics, García-Bernete et al. (2024) recently studied the relation between the PAH feature at 6.2 and the S(1) line to establish alternative ways to determine the presence of an AGN in the absence of high IP lines. The authors suggest that ratios of 6.2/S(1) larger than 12 are dominated by SF processes, whereas ratios below this value reflect an increase on the due to the heating of shocks/AGN. Stronger radiation fields would further destroy the PAHs, reducing the ratio below 5. For CGCG 007-025, we report a lower limit of , which would fall in the AGN demarcation. Nevertheless, this trend has been established in (super-)solar metallicity environments as well as in confirmed AGN host galaxies.At lower metallicities, it is still to be determined whether this ratio can be attributed to the harder ionizing spectrum (which would destroy the PAH complexes and increase the S(1) flux, simultaneously), or whether they are shaped simply because of the lack of PAH molecules in the first place.
@@
It is worth noting that the PAH emission peaks at the location of high-IP lines ([O IV] and [Ne V]), with no significant PAH detections in regions outside of these high-IP boundaries. In the scenario of PAH destruction in the presence of a highly ionizing field, this spatial anti-correlation seems contradictory: smaller PAHs should survive preferentially in lower-ionization environments. An alternative to PAH destruction is in-situ formation, where the source producing [O IV] and [Ne V]emission could also be the origin of PAH molecules. For example, the formation of PAH has previously been seen in the dusty winds of massive and/or Wolf-Rayet (WR) stars (Lau et al., 2022; Taniguchi et al., 2025). Although WR stars seem capable of producing [Ne V] emission (see Tarantino et al., 2024; Hunt et al., 2025), current photoionization models of massive stars cannot reproduce the mid-IR line ratios seen in CGCG 007-025. Further theoretical and observational efforts are thus required to understand the production of high-ionization photons and the formation of dust and complex molecules in low-metallicity starburst environments.
6 Conclusions
@@
The 11.3 PAH feature is significantly detected at spaxel level (left panel of Figure 2). We explore the extension and location of this feature. The emission is spatially confined to a region of pc in the vicinity of the brightest clumps on the galaxy.
@@
We do not find evidence of any other PAH feature in the integrated spectrum of CGCG 007-025. We place an upper limit detection on the 6.2 PAH complex of .
The exclusive detection of the 11.3 feature indicates a PAH population dominated by larger, neutral molecules. Moreover, the 6.2/11.3 ratio, in conjunction with a large [Ne III]/[Ne II] ratio (Figure 4), hint to the presence of a strong ionizing source, with [Ne III]/[Ne II] ratios compatible with those seen in harsh environments like AGNs or other BCDs.
@@ Appendix A Data reduction
The data were reduced using the JWST data reduction pipeline version 1.17.1.dev with the CRDS 1338 context. Briefly, in Stage 1 corrections at detector level are applied to produce data in count rate units. These include corrections for dark current, masking of detector cosmetics, and removal of strong jumps introduced by the impact of cosmic rays. In Stage 2, data is flat-field corrected, and wavelength and flux calibrated. A first inspection of the count rate images revealed intense cosmic ray events all across the detector which are difficult to mask. At the time of reducing the data, the developer version of the pipeline was the only release which included the straylight correction step, a crucial step for correcting these intense cosmic ray events. Background subtraction is performed in Stage 3 using a master sky background created with the dedicated sky observations of each object. We opted for this option rather than a pixel-by-pixel correction, since the later introduces more noise in the final data products. Stage 3 also combines the individual exposures into a single one and reconstructs the 3D datacubes per band. We also enabled a residual fringe correction in Stage 2 to account for any residual fringing not corrected during Stage 1. For extended sources, the simple fringe-flat correction applied in Stage 1 is usually enough to reduce its impact on the data products. However, the galaxy displays a PSF-like structure, due to an unresolved compact emission, making the fringe patterns vary systematically across the PSF and requiring further corrections to create the final extended mid-IR emission datacubes. We chose the emsm algorithm as a first step to minimize the wiggle patterns of the PSF. The additional PSF-like emission modeling and subtraction was done following the Mingozzi et al. (2025) procedure. Besides our best efforts to correct from fringing and PSF undersampling, some spaxels situated at the peak of the mid-IR emission are still affected by these wiggling patterns.
Appendix B Additional material
@@
In Table 1 we compile the complete list of emission lines included in pahfit for the modeling of the PAH emitting region. Table 2 displays the unattenuated emission line measurements from pahfit together with the flux of the 11.3 m PAH feature and the upper limit on the 6.2 m PAH feature.
| Fine structure lines | H I lines | lines | ||||
| Line name | IP | Transition | Line name | |||
| [eV] | [] | [] | [] | |||
| [Fe II] | 7.9 | 5.34 | Hu(10-6) | 5.129 | S(8) | 5.053 |
| [Ar II] | 15.76 | 6.99 | Hu(9-6) | 5.908 | S(7) | 5.511 |
| [Ar III] | 27.63 | 8.99 | Pf(6-5) | 7.45985 | S(6) | 6.109 |
| [S IV] | 34.83 | 10.51 | Hu(8-6) | 7.503 | S(5) | 6.909 |
| [Ne II]⋆ | 21.56 | 12.81 | 10-7 | 8.76 | S(4) | 8.026 |
| [Ne V]⋆ | 97.12 | 14.32 | 9-7 | 11.3088 | S(3) | 9.665 |
| [Ne III]⋆ | 40.96 | 15.55 | Hu(7-6) | 12.37188 | S(2) | 12.278 |
| [S III] | 23.34 | 18.71 | 8-7 | 19.062 | S(1) | 17.035 |
| [O IV]⋆ | 54.95 | 25.89 | ||||
| [Fe II] | 7.9 | 25.99 | ||||
| Fine structure lines | lines | PAH | |||
| Line name | Flux | Line name | Flux | Feature | Flux |
| [Fe II]5.34 | 0.4 0.3 | S(1) | 0.80 0.03 | 6.2 | |
| [Ar II]6.99 | 0.29 0.08 | 11.3 | 5.1 0.4 | ||
| [Ar III]8.99 | 3.86 0.08 | ||||
| [S IV]10.51 | 31.2 0.4 | ||||
| [Ne II]12.81 | 2.22 0.03 | ||||
| [Ne V]14.32 | 0.38 0.02 | ||||
| [Ne III]15.55 | 35.3 0.4 | ||||
| [S III]18.71 | 43.4 0.4 | ||||
| [O IV]25.89 | 24.4 0.4 | ||||
| [Fe II]25.99 | 1.2 0.3 | ||||
In Figure 6 we show the SFR-PAH relation for the 11.3 (left) and 6.2 (right) PAH features. The BCDs sample from Xie & Ho (2019) is displayed as purple diamonds, with open symbols representing upper limits on the PAH luminosities. The different SFR-PAH calibrations from Xie & Ho (2019) are shown in purple, with other calibrations from the literature displayed in green (Mordini et al., 2021). The measurements from CGCG 007-025 are shown in orange.

### v2

@@ V.1 What PAH and line ratios reveal about CGCG 007-025
Complementing these diagnostics, García-Bernete et al. (2024) recently studied the relation between the PAH feature at 6.2 and the S(1) line to establish alternative ways to determine the presence of an AGN in the absence of high IP lines. The authors suggest that ratios of 6.2/S(1) larger than 12 are dominated by SF processes, whereas ratios below this value reflect an increase on the due to the heating of shocks/AGN. Stronger radiation fields would further destroy the PAHs, reducing the ratio below 5. For CGCG 007-025, we report a upper limit of , which would fall in the AGN demarcation. Nevertheless, this trend has been established in (super-)solar metallicity environments as well as in confirmed AGN host galaxies. At lower metallicities, it is still to be determined whether this ratio can be attributed to the harder ionizing spectrum (which would destroy the PAH complexes and increase the S(1) flux, simultaneously), or whether they are shaped simply because of the lack of PAH molecules in the first place.
@@
It is worth noting that the PAH emission peaks at the location of high-IP lines ([O IV] and [Ne V]), with no significant PAH detections in regions outside of these high-IP boundaries. In the scenario of PAH destruction in the presence of a highly ionizing field, this spatial anti-correlation seems contradictory: smaller PAHs should survive preferentially in lower-ionization environments. An alternative to PAH destruction is in-situ formation, where the source producing [O IV] and [Ne V] emission could also be the origin of PAH molecules. For example, the formation of PAH has previously been seen in the dusty winds of massive and/or Wolf-Rayet (WR) stars (Lau et al., 2022; Taniguchi et al., 2025). Although WR stars seem capable of producing [Ne V] emission (see Tarantino et al., 2024; Hunt et al., 2025a), current photoionization models of massive stars cannot reproduce the mid-IR line ratios seen in CGCG 007-025. Further theoretical and observational efforts are thus required to understand the production of high-ionization photons and the formation of dust and complex molecules in low-metallicity starburst environments.
VI Conclusions
@@
The 11.3 PAH feature is significantly detected at spaxel level (left panel of Figure 2). We explore the extension and location of this feature. The emission is spatially confined to a region of pc, offset from the brightest clumps on the galaxy.
@@
We find evidence of another PAH feature at 12.7 , but the usually bright features at 6.2, 7.7 or 8.6 are absent in CGCG 007-025.
The detection of the 11.3 and 12.7 feature may indicate a PAH population dominated by larger, neutral molecules. Moreover, the 6.2/11.3 ratio, in conjunction with a large [Ne III]/[Ne II] ratio (Figure 3), hint to the presence of a strong ionizing source, with [Ne III]/[Ne II] ratios compatible with those seen in harsh environments like AGNs or other BCDs.
@@ Appendix A Data reduction
The data were reduced using the JWST data reduction pipeline version 1.17.1.dev with the CRDS 1338 context. Briefly, in Stage 1 corrections at detector level are applied to produce data in count rate units. These include corrections for dark current, masking of detector cosmetics, and removal of strong jumps introduced by the impact of cosmic rays. In Stage 2, data is flat-field corrected, and wavelength and flux calibrated. A first inspection of the count rate images revealed intense cosmic ray events all across the detector which are difficult to mask. At the time of reducing the data, the developer version of the pipeline was the only release which included the straylight correction step, a crucial step for correcting these intense cosmic ray events. Background subtraction is performed in Stage 3 using a master sky background created with the dedicated sky observations. We opted for this option rather than a pixel-by-pixel correction, since the later introduces more noise in the final data products. Stage 3 also combines the individual exposures into a single one and reconstructs the 3D datacubes per band. We also enabled a residual fringe correction in Stage 2 to account for any residual fringing not corrected during Stage 1. For extended sources, the simple fringe-flat correction applied in Stage 1 is usually enough to reduce its impact on the data products. However, the galaxy displays a PSF-like structure, due to an unresolved compact emission, making the fringe patterns vary systematically across the PSF and requiring further corrections to create the final extended mid-IR emission datacubes. We chose the emsm algorithm as a first step to minimize the wiggle patterns of the PSF. The additional PSF-like emission modeling and subtraction was done following the Mingozzi et al. (2025) procedure. Besides our best efforts to correct from fringing and PSF undersampling, some spaxels situated at the peak of the mid-IR emission are still affected by these wiggling patterns. We also verified that the astrometry of the MIRI/MRS datacubes matched the World Coordinate System (WCS) of the available HST data of this galaxy, as well as the WCS of the MUSE dataset from which we obtained the H contours.
Appendix B Absence of typical PAH features in CGCG 007-025
Despite of the bump at 11.3 m, the integrated spectrum of CGCG 007-025 does not display clear evidence of other PAH features. In Figure 5, we compare the integrated spectrum of CGCG 007-025 (gray line) with the PDRs4All template for HII regions (pink line, Chown et al., 2024). Both spectra have been continuum subtracted, and the template is also re-scale to match the flux of the 11.3 m feature. The spectrum of CGCG 007-025 shows no excess of emission at 6.2 and 7.7 m (left panel), indicating no evidence of these features in the aperture. However, the spectrum displays a slight excess of emission at 12.7 m, matching the aforementioned template (right panel). We measure the flux associated with the 12.7 feature on the continuum subtracted spectrum using a Drude profile. The integrated flux, corrected for attenuation, can be found in Table 1. Following Lai et al. (2025), we derive upper limits on the undetected PAHs by estimating the integrated fluxes of the 6.2, 7.7 and 8.6 bands. We define as the product of the width of the PAH profile and rms of the residuals from PAHFIT in the wavelength range 9.68–10.43 , where no PAH or emission lines are observed. The width of the PAH bands are defined as the interval spanning one FWHM, centered on the central wavelength (Smith et al., 2007). For the PAH complexes 6.2 and 7.7, the wavelength range expands from half a FWHM below the bluest subcomponent to half a FWHM above the reddest subcomponent. The upper limits on the PAH bands can be found on Table 1.
Appendix C Additional material
@@
In Table 1 we compile the well-detected list of emission lines included in pahfit for the modeling of the PAH emitting region, as well as the unattenuated emission line measurements from pahfit. We also included the flux of the 11.3 and 12.7 m PAH features and the upper limits on the 6.2, 7.7 and 8.6 m PAH features.
| Fine structure lines | lines | PAH | ||||||
| Line name | IP | Flux | Line name | Flux | Feature | Flux | ||
| [eV] | [] | [] | [] | [] | [] | |||
| [Fe II] | 7.9 | 5.34 | 0.4 0.3 | S(1) | 17.035 | 0.80 0.03 | 6.2 | |
| (0.5 0.3) | (1.00 0.03) | () | ||||||
| [Ar II] | 15.76 | 6.99 | 0.29 0.08 | 7.7 | ||||
| (0.3 0.1) | () | |||||||
| [Ar III] | 27.63 | 8.99 | 3.86 0.08 | 8.6 | ||||
| (5.8 0.1) | () | |||||||
| [S IV] | 34.83 | 10.51 | 31.2 0.4 | 11.3 | 3.9 0.4 | |||
| (49.0 0.6) | (5.1 0.4) | |||||||
| [Ne II]⋆ | 21.56 | 12.81 | 2.22 0.03 | 12.7 | 7.0 0.4 | |||
| (2.38 0.03) | (7.5 0.4) | |||||||
| [Ne V]⋆ | 97.12 | 14.32 | 0.38 0.02 | |||||
| (0.41 0.02) | ||||||||
| [Ne III]⋆ | 40.96 | 15.55 | 35.3 0.4 | |||||
| (40.3 0.5) | ||||||||
| [S III] | 23.34 | 18.71 | 43.4 0.4 | |||||
| (55.1 0.5) | ||||||||
| [O IV]⋆ | 54.95 | 25.89 | 24.4 0.4 | |||||
| (25.4 0.4) | ||||||||
| [Fe II] | 7.9 | 25.99 | 1.2 0.3 | |||||
| (1.2 0.3) | ||||||||